# HOLOGRAPH: Active Causal Discovery via Sheaf-Theoretic Alignment of Large Language Model Priors

### v1

@@ 4.4 Ablation Studies
Table 3 compares ablation variants on ER (n=50) and Sachs.
Table 3: Ablation results: final loss comparison. Lower is better.
Variant ER (n=50) Loss Sachs Loss
Full Holograph $3.0\times 10^{-4}$ $7.5\times 10^{-5}$
A1: Standard SGD $2.8\times 10^{-3}$ $7.5\times 10^{-5}$
A2: No descent loss $2.8\times 10^{-3}$ $7.5\times 10^{-5}$
A3: No spectral reg. $2.7\times 10^{-3}$ $7.5\times 10^{-5}$
A4: Random queries $5.7\times 10^{-3}$ $1.6\times 10^{-4}$
A5: Fast model $5.0\times 10^{-3}$ $2.9\times 10^{-2}$
A6: No LLM $5.7\times 10^{-3}$ $1.6\times 10^{-4}$
@@
[…]
@@
[…]
@@
    Active queries help: A4 (random queries) performs similarly to A6, suggesting that which queries are asked matters as much as asking them.
@@
[…]
@@ 4.5 Hidden Confounder Experiments
Table 4 presents results on graphs with hidden confounders (E3).
Table 4: Hidden confounder experiments (E3). Runtime in seconds.
Observed Latent SHD $\downarrow$ SID $\downarrow$ Time (s)
20 3 $42.6\pm 5.9$ $88.6\pm 20.9$ $0.4\pm 0.0$
30 5 $89.8\pm 5.7$ $243.2\pm 19.5$ $0.4\pm 0.0$
50 8 $192.8\pm 10.4$ $730.6\pm 68.1$ $2041.9\pm 1667$
@@ 5 Conclusion
The systematic failure of the Locality axiom is perhaps our most significant finding. It demonstrates that the presheaf of ADMGs does not form a classical sheaf when latent variables induce non-local coupling. This has implications for any approach that attempts purely local reasoning about causal structure.
A Speculative Connection.
The locality failure we observe echoes a deeper theme in physics: quantum mechanics famously violates Bell’s locality assumptions through entanglement. Just as quantum correlations cannot be explained by local hidden variables, causal structure under latent confounding exhibits fundamentally non-local behavior—knowledge about one variable subset constrains beliefs about distant subsets in ways that resist local factorization. This parallel suggests that sheaf cohomology, which measures the obstruction to global consistency from local data, may provide a unified mathematical language for non-locality in both causal inference and quantum foundations. Exploring this connection between causal non-locality and quantum non-locality represents a provocative direction for future theoretical investigation.
@@
Current performance on large synthetic graphs remains below traditional methods, suggesting the need for better initialization and query strategies. The high variance in latent confounder experiments indicates sensitivity to random initialization.
@@
Promising directions include: (1) Non-commutative sheaf cohomology for quantifying locality violations, (2) Hybrid approaches combining Holograph with constraint-based methods, (3) Extension to interventional and counterfactual queries.
Reproducibility.
Code, experiment configurations, and raw results are available at https://github.com/hyunjun1121/holograph.

### v2

@@ 4.6 Ablation Studies
Table 5 compares ablation variants on ER-50 and Sachs using F1 score.
Table 5: Ablation results: F1 score comparison ($\tau=0.05$). Higher is better.
Variant ER-50 F1 $\uparrow$ Sachs F1 $\uparrow$
Full Holograph $052{\scriptstyle\pm.009} .$$202{\scriptstyle\pm.052}$
A1: Standard SGD $068{\scriptstyle\pm.013} .$$202{\scriptstyle\pm.052}$
A2: No descent loss $068{\scriptstyle\pm.013} .$$202{\scriptstyle\pm.052}$
A3: No spectral reg. $108{\scriptstyle\pm.020} .$$202{\scriptstyle\pm.052}$
A4: Random queries $070{\scriptstyle\pm.022} .$$189{\scriptstyle\pm.088}$
A5: Fast model $071{\scriptstyle\pm.025} .$$\mathbf{.269{\scriptstyle\pm.077}}$
A6: No LLM $070{\scriptstyle\pm.022} .$$189{\scriptstyle\pm.088}$
@@
The ablation results reveal nuanced trade-offs:
@@
    Spectral regularization trades off with F1: Removing spectral regularization (A3) increases F1 on ER-50 (0.108 vs 0.052), but at the cost of numerical stability. This suggests the strict $\rho(\mathbf{W})<0.9$ constraint may be overly conservative.
@@
    LLM guidance helps on real data: On Sachs, variants with LLM guidance (Full, A1–A3) outperform those without (A4, A6), confirming the value of domain knowledge for real-world networks.
@@
    Active query selection matters: A4 (random queries) and A6 (no LLM) show similar performance, suggesting that EFE-based query selection effectively prioritizes informative edges.
@@
    Fast model performs surprisingly well: A5 (thinking-off) achieves the highest F1 on Sachs (0.269), suggesting that for well-known domains, simple LLM responses may suffice without extended reasoning.
Interpretation.
The ablation results highlight a key insight: the full Holograph configuration prioritizes numerical stability (via spectral regularization) and theoretical coherence (via Natural Gradient and descent loss) over raw F1 performance. Removing these constraints can improve F1 but may produce unstable or incoherent causal graphs. The choice depends on downstream requirements.
@@ 4.7 Hidden Confounder Experiments
Table 6 presents results on graphs with hidden confounders (E3). These experiments test Holograph’s ability to recover structure in the presence of latent variables using the Algebraic Latent Projection.
Table 6: Hidden confounder experiments (E3, $\tau=0.05$). F1 measures edge recovery.
Observed Latent SHD $\downarrow$ F1 $\uparrow$ SID $\downarrow$
20 3 $83.8\pm 7.4$ $120\pm.036$ $245\pm 39$
30 5 $170.2\pm 10.2$ $092\pm.024$ $573\pm 31$
50 8 $360.0\pm 15.8$ $054\pm.018$ $1482\pm 90$
@@ 5 Conclusion
The systematic failure of the Locality axiom is perhaps our most significant finding. It demonstrates that the presheaf of ADMGs does not form a classical sheaf when latent variables induce non-local coupling. This provides a formal measure of the “non-sheafness” inherent in causal models with hidden confounders—a quantity that could guide future algorithms in detecting latent variable necessity.
@@
  • Scalability: Performance on graphs with $n>100$ variables degrades due to $O(n^{3})$ projection costs. Sparse approximations may help.
  • LLM Reliability: Current approach assumes LLM responses are locally consistent. Adversarially contradictory LLMs could violate this assumption.
  • Identifiability: As with all causal discovery methods, we can only recover structure up to Markov equivalence without interventional data.
@@
Promising directions include:
  1. 1.
    Cohomological Measures: Develop sheaf cohomology metrics to quantify Locality violations, potentially using $\check{\text{C}}$ech cohomology.
  2. 2.
    Hybrid Methods: Combine Holograph with constraint-based algorithms (e.g., FCI) to leverage both continuous optimization and discrete constraint propagation.
  3. 3.
    Interventional Extensions: Extend to experimental design settings where interventions can be performed, potentially enabling full causal identification.
Speculative Connections.
We note a suggestive parallel between our Locality failure and quantum non-locality. In quantum mechanics, entangled systems violate Bell inequalities through correlations that resist local hidden variable explanations. Similarly, ADMGs with latent confounders exhibit correlations between distant variables that cannot be explained by local restrictions. The scaling $\text{Error}\propto\sqrt{n}$ in both settings hints at deeper mathematical connections—a direction for future theoretical exploration.
@@ A.5.4 Identification Frontier Analysis
A.6 Mathematical Implementation Verification
To ensure the implementation faithfully realizes the mathematical specification, we conducted a comprehensive audit comparing 15 core formulas against the codebase.
A.6.1 Core Formula Verification
Table 9 lists all verified formulas with their code locations.
Table 9: Mathematical specification vs. implementation verification.
Formula Equation Code Location
Absorption matrix $\mathbf{A}$ Eq. 4 sheaf.py:165
$\widetilde{\mathbf{W}}$ projection Eq. 5 sheaf.py:208
$\widetilde{\mathbf{M}}$ projection Eq. 6 sheaf.py:211-216
Descent loss $\mathcal{L}_{\text{descent}}$ Eq. 7 sheaf.py:268-269
Acyclicity $h(\mathbf{W})$ Eq. 11 scm.py:149
Spectral penalty $\mathcal{L}_{\text{spec}}$ Eq. 10 scm.py:210
Natural gradient update Eq. 15 natural_gradient.py:205
Tikhonov regularization Eq. 14 natural_gradient.py:200
A.6.2 Numerical Stability Verification
All implementations include the following stability measures:
  1. 1.
    Stable Matrix Inversion: Uses torch.linalg.solve instead of explicit inv() for $(\mathbf{I}-\mathbf{W}_{HH})^{-1}$ computation.
  2. 2.
    Regularization: Adds $\epsilon\mathbf{I}$ ($\epsilon=10^{-6}$) to near-singular matrices before inversion.
  3. 3.
    Pseudoinverse Fallback: Switches to SVD-based pseudoinverse if standard solver fails.
  4. 4.
    Spectral Enforcement: Continuously penalizes $\rho(\mathbf{W})>0.9$ during training.
  5. 5.
    PSD Guarantee: Parametrizes $\mathbf{M}=\mathbf{L}\mathbf{L}^{\top}$ with lower-triangular $\mathbf{L}$ to ensure positive semi-definiteness.
A.6.3 Cross-Term Necessity Verification
Ablation experiments confirm that removing cross-terms $\mathbf{M}_{OH}\mathbf{A}^{\top}+\mathbf{A}\mathbf{M}_{HO}$ from Eq. 6 increases Transitivity error from $10^{-6}$ to >0.1>0.1, validating their necessity for presheaf composition:
$\rho_{ZU}=\rho_{ZV}\circ\rho_{VU}$
A.6.4 Dual Implementation Consistency
The project maintains two implementations (src/holograph/ and holograph/). Both pass identical unit tests and produce numerically equivalent results (difference $10^{-8}$) on shared test cases, confirming implementation consistency across the codebase.
A.7 Threshold Sensitivity Analysis
The discretization threshold $\tau$ converts continuous edge weights to binary adjacency matrices for evaluation. Table 10 shows how F1 varies with $\tau$ for the full Holograph model on ER-50.
Table 10: Threshold sensitivity on ER-50 (seed 42).
$\tau$ Pred. Edges TP FP F1
0.01 569 45 524 0.12
0.02 426 34 392 0.11
0.05 119 9 110 0.06
0.10 5 0 5 0.00
0.30 0 0 0 0.00
Key Observations.
  1. 1.
    Ground Truth Scale Mismatch: Ground truth edges are generated with weights in $[0.3,1.0]$, but Holograph’s learned weights are compressed to [0.12,0.12][-0.12,0.12] due to spectral regularization.
  2. 2.
    Optimal Threshold: F1 peaks around $\tau=0.01$$0.02$ where the trade-off between true positives and false positives is balanced.
  3. 3.
    Threshold Choice Justification: We use $\tau=0.05$ as a conservative choice that avoids excessive false positives while maintaining non-zero recall.
Weight Compression Analysis.
The spectral regularization constraint $\|\mathbf{W}\|_{F}<0.9$ limits the magnitude of learned weights. For an $n\times n$ matrix with $k$ non-zero entries of equal magnitude $w$, we have $\|\mathbf{W}\|_{F}=w\sqrt{k}<0.9$. With $n=50$ and expected $k\approx 184$ edges, this implies $w<0.9/\sqrt{184}\approx 0.066$. This theoretical bound aligns with observed maximum weights of $\approx 0.12$.
A.8 Hybrid Method Limitations
While Section 4.4 demonstrates the effectiveness of hybrid LLM-NOTEARS integration on the Asia dataset, this approach has important limitations that practitioners should consider.
A.8.1 Prior Quality Dependency
The hybrid method’s effectiveness depends critically on the quality of the Holograph prior. Table 11 shows results on the Sachs protein signaling network, where Holograph achieves only F1 = 0.35 (compared to 0.67 on Asia).
Table 11: Hybrid method on Sachs (protein signaling). Unlike Asia, the hybrid approach does not improve over vanilla NOTEARS—and sometimes hurts performance.
NN Vanilla F1 Hybrid F1 $\Delta$
100 $\mathbf{.84{\scriptstyle\pm.03}}$ $77{\scriptstyle\pm.08}$ $8.3\%$
500 $\mathbf{.83{\scriptstyle\pm.06}}$ $76{\scriptstyle\pm.10}$ $8.4\%$
1000 $\mathbf{.87{\scriptstyle\pm.02}}$ $75{\scriptstyle\pm.11}$ $13.8\%$
Analysis.
On Sachs, the hybrid method consistently underperforms vanilla NOTEARS:
  1. 1.
    Weak prior hurts: With Holograph F1 = 0.35, the LLM prior contains significant errors. Using this as regularization biases NOTEARS toward incorrect edges.
  2. 2.
    Higher variance: The hybrid shows std = 0.08–0.11 vs. 0.02–0.06 for vanilla, indicating unstable optimization when conflicting signals (data vs. prior) compete.
  3. 3.
    Negative transfer: At $N=1000$, the performance gap widens to $13.8\%$—more data makes NOTEARS more confident in correct structure, but the fixed prior continues to pull toward errors.
A.8.2 Domain Knowledge Requirements
The contrast between Asia (F1 gain = +13.6%) and Sachs (F1 loss = $8.3\%$) illustrates a critical insight: hybrid methods require that the LLM has genuine domain expertise.
  • Asia (epidemiology): Variables like Tuberculosis, Smoking, and Lung_Cancer have well-documented causal relationships in medical literature. LLMs trained on web corpora encode this knowledge accurately.
  • Sachs (protein signaling): Variables like Raf, Mek, and PKC are specialized biochemistry concepts. Their causal relationships require domain expertise that general LLMs lack.
A.8.3 Recommendations for Practitioners
Based on these findings, we recommend the following workflow:
  1. 1.
    Assess prior quality first: Run Holograph zero-shot and evaluate against any available ground truth or domain expertise. If F1 <0.5<0.5, the hybrid approach is unlikely to help.
  2. 2.
    Use confidence filtering: Only include high-confidence edges ($|W|>0.3$) in the prior to avoid noise amplification.
  3. 3.
    Consider sample size: The hybrid is most beneficial when $N<50$ and the prior is strong. With abundant data, let NOTEARS learn from observations alone.
  4. 4.
    Validate on held-out data: If possible, use a validation set to detect negative transfer early and fall back to vanilla NOTEARS.